OpenStudy (anonymous):

How do I find the radius of a circle with a center at (-1, -2), which also passes through point (0,2)?

OpenStudy (amoodarya):

\[(x-x_c)^2+(y-y_c)^2=r^2\\center=(-1,-2)\\so\\(x+1)^2+(y+2)^2=r^2\\\] the circle passes through (0,2)  so check the point in equation to find r can you go on ?

OpenStudy (anonymous):

@amoodarya I'm still a little confused. Could you explain further?

OpenStudy (anonymous):

The previous post used the standard form of a circle at center (h ,k) Let h=xc let k= yc then you have a point (0,2) (0,2) (x+1)^2 + (y+2)^2 = r^2 (0+1)^2 + (2+2)^2 = r^2 1+4 = r^2 5= r^2 \[r=\sqrt{5}\]

OpenStudy (anonymous):

Ok. Thanks